Dan is on a 3-day hike. The first day he hikes 1/3 of the trail. The next day he hikes 1/3 of the trail plus 2.4 kilometers. The last day he hikes 5.6 kilometers to the end of the trail. How long is the trail?
Answer by rapaljer(4671)

Let x = length of trail

The sum of the first day + second day + third day = Total Trip = x
(1/3)x + (1/3)x +2.4 + 5.6 =x

(2/3)x + 8 = x

Multiply both sides by 3:
2x + 24 = 3x

Subtract 2x from each side:
24 = 3x-2x
24 km =x

R^2